# Robust iterative methods for linear systems with saddle point structure

### v1

@@ 3.1 The symmetric case
In the symmetric case, we have $B=C$ and $A=A^{T}$, and we compute a sparse approximate nullspace basis, $\widetilde{Z}$, of $B^{T}$. We then assume, see Section 2 for conditions when ${\mathcal{N}}^{s}=\widetilde{Z}^{T}A\widetilde{Z}$ is positive definite. Since the new scheme involves sparse approximations and just used as a preconditioner, in the following numerical experiments, we assume that $\widetilde{\mathcal{N}}^{s}$ is positive definite, even if $A$ is indefinite or semidefinite. This allows us to cover a larger number applications than in our initial assumption. The pseudocode of the new scheme for the symmetric case is given in Algorithm 5.
Optionally, we can further $M$-orthogonalize the columns of $\widetilde{Z}$ with respect to $M=A$ using the modified Gram-Schmidt (MGS) procedure given in Algorithm 4. Since it may be too expensive to orthogonalize against all vectors, we also adopt a ’partial’ orthogonalization against a certain number of previous vectors, which we refer to as the window size $w$ (used in line 4) and we employ numerical dropping with a threshold $\tau$ (used in line 7) to ensure sparsity. Although $M$-orthogonalization can be achieved during the oblique conjugation process in Algorithm 2 for the last $(n-k)$ columns of $V$, we prefer to keep it as a separate stage in our new scheme, since it requires the matrix $M=A$ which may change when solving multiple linear systems with different right-hand sides even though $B$ remains the same.
@@
1:$Z\in\mathbb{R}^{n\times(n-k)}$, $M\in\mathbb{R}^{n\times n}$triangleright $z_{i}$: is the $i^{th}$ column of $Z$
2:$\bar{Z}\in\mathbb{R}^{n\times(n-k)}$triangleright $\bar{z}_{i}$: is the $i^{th}$ column of $\bar{Z}$
3:$\bar{z}_{1}=z_{1}/\sqrt{z_{1}^{T}Mz_{1}}$
@@
1:$A\in\mathbb{R}^{n\times n}$ with $A=A^{T}$, $B\in\mathbb{R}^{n\times m}$, $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ with $m<n$
2:$x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$
@@
  • Find a particular solution, ,$\hat{z}_{1}\in\mathbb{R}^{n}$, of $-B^{T}z_{1}=t_{2}$ via LSQR
  • Solve $\widetilde{W}^{T}\widetilde{Z}^{T}A\widetilde{Z}\widetilde{W}(\widetilde{W}^{-1}u)=\widetilde{W}^{T}\widetilde{Z}^{T}(t_{1}-A\hat{z_{1}})$ via CG
@@ 3.2 Generalized case
1:$A\in\mathbb{R}^{n\times n}$ with $A\neq A^{T}$, $B\in\mathbb{R}^{n\times m}$, $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ with $m<n$
2:$x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$
@@
  • Find a particular solution, $\hat{z}_{1}\in\mathbb{R}^{n}$, of $-B^{T}z_{1}=t_{2}$ via LSQR
  • Solve $\widetilde{W}^{T}\widetilde{Z}^{T}A\widetilde{Z}\widetilde{W}(\widetilde{W}^{-1}u)=\widetilde{W}^{T}\widetilde{Z}^{T}(t_{1}-A\hat{z_{1}})$ via fGMRES with the preconditioner $N=I+\widetilde{W}^{T}\mathcal{N}^{J}\widetilde{W}$ (systems involving the preconditioner are solved via MRS), where $\mathcal{N}^{J}=\widetilde{Z}^{T}(\frac{A-A^{T}}{2})\widetilde{Z}$
@@ 3.3 General case
1:$A\in\mathbb{R}^{n\times n}$ with $A\neq A^{T}$, $B,C\in\mathbb{R}^{n\times m}$ with $B\neq C$, $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ with $m<n$
2:$x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$
@@
  • Find a particular solution, ,$\hat{z}_{1}\in\mathbb{R}^{n}$, of $-C^{T}z_{1}=t_{2}$ via LSQR
  • Solve $\widetilde{W}^{T}\widetilde{Z}^{T}A\widetilde{U}\widetilde{W}(\widetilde{W}^{-1}u)=\widetilde{W}^{T}\widetilde{Z}^{T}(t_{1}-A\hat{z_{1}})$ via fGMRES with a preconditioner $N=I+\widetilde{W}^{T}\widetilde{\mathcal{N}}^{J}\widetilde{W}$ (systems involving the preconditioner are solved via MRS) where $\widetilde{\mathcal{N}}^{J}=(\widetilde{Z}^{T}A\widetilde{U}-\widetilde{U}^{T}A^{T}\widetilde{Z})/2$
@@ 4 Test problems
For the generalized case, eight matrices are used (given in Table 2). The first three, namely rajat04, rajat14 and fpga_trans_01, are from the SuiteSparse Matrix Collection. The remaining five matrices arise in the steady-state solution of a linearized 2D Navier-Stokes equations and are obtained from IFISS software [10] for the lid-driven cavity problem with a spatial discretization based on $Q_{2}/Q_{1}$ elements and a uniform grid, varying the Reynolds number between $100$ and $900$. We note that, in this case, the $B$ matrix generated by the IFISS software is numerically rank-deficient. Therefore, in the new scheme, we compute a nullspace basis of size $m-n+1$. In Table 2, we provide the smallest and largest eigenvalues of $A+A^{T}$. For three of the test problems in this set, $A+A^{T}$ is positive definite but close to singular, while for the remaining five problems it is robustly positive definite.
[…]
@@
1 reorientation_1 396 281 7,326 $-1.7\times 10^{6}$ $1.0\times 10^{9}$
@@
6 goddardRocketProblem_2 434 433 9,058 $-1.1\times 10^{3}$ $1.1\times 10^{3}$
@@
6 drivencavity_Re500 578 81 10,814 $-8.2\times 10^{-4}$ 1.0
7 drivencavity_Re700 578 81 10,814 $-2.1\times 10^{-3}$ 1.0
8 drivencavity_Re900 578 81 10,814 $-9.1\times 10^{-3}$ 1.0
@@
8 random2 (cond.$2.89\times 10^{9}$ ) 100 90 554
9 random3 (cond.$2.26\times 10^{4}$ ) 1,000 900 3,079
@@ 5 Experimental framework
The following numerical results that we present were obtained using Matlab R2023b on an Apple Macbook Air with an M2 processor and 8 GB of memory. As a baseline of comparison, we use restarted GMRES(m) with $m=10$, preconditioned with MATLAB’s incomplete $LU$ method with pivoting and thresholding, ILUTP($\tau$), where $\tau$ is the drop tolerance. We apply reverse Cuthill-McKee reordering to the systems before computing the incomplete factors by ILUTP($\tau$) to improve the robustness of the baseline method, see [5]. Here, GMRES(m) is preconditioned from the left and right. We note that for the symmetric case, an incomplete $LDL^{T}$ factorization based preconditioner would be more effective and suitable as a baseline, see [15]. Unfortunately, it is not available as a built-in function in MATLAB, and we were unable to compile the source code for the Apple M processor architecture. Therefore, we use ILUTP($\tau$) for all cases, including the symmetric case. For our new schemes, we use the restarted and right preconditioned flexible GMRES; fGMRES(m), with $m=10$, as the solver for the outer iteration. We chose fGMRES(m) because the proposed preconditioners (comprised of other iterative solvers) inherently change between iterations. The new schemes have a number of parameters: (a) drop tolerances and thresholds: $\tau_{SAROC},\rho_{SAROC},\tau_{FSAI},\rho_{FSAI}$ and (b) Stopping tolerances for the outer fGMRES, inner LSQR, inner fGMRES and innermost MRS (or inner GC). Furthermore, if $M$-orthogonalization is performed on the nullspace basis, we include the window size ($w_{MGS}$) and drop tolerance ($\tau_{MGS}$) parameters. To illustrate the robustness of the new solver class, we experiment with small, large, and mixed thresholds and tolerances. These are given in Table 4. For ILUTP($\tau$), $\tau$ is set to $10^{-5},10^{-4},10^{-3},10^{-2}$ and $10^{-1}$, although drop tolerances of $10^{-3}-10^{-1}$ are preferred more often in practice. One of the main difficulties for classical incomplete factorization-based preconditioners is that there is no easy way to predict a good drop tolerance that would work well for a given problem. Sometimes, it is even counterintuitive that decreasing the drop tolerance might result in encountering a zero pivot.
@@
For both the new schemes and the baseline, we start the iterations with the zero vector and stop when the $L_{2}$ norm of the relative residual drops below $10^{-5}$ or the maximum number ($1,000$) of iterations is reached. Note that for the case that good starting vectors are available (for example if one has a sequence of slightly changing matrices or right hand sides) then the convergence will be substantially faster.
@@
The residual norms are obtained as a by-product of fGMRES and GMRES. While right preconditioned fGMRES produces the true residual norm, and GMRES produces the preconditioned residual norm due to ILUTP-based left and right preconditioning. For these reasons, we also calculate and report the final true relative residual norms. For the new schemes, for the symmetric case, the inner LSQR and CG iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and $\epsilon_{innermost}$ is not used. For the generalized and general case, inner LSQR and fGMRES iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and the innermost MRS iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{innermost}$. Since our new schemes are based on multiple nested loops and external functions, a fair comparison in terms of the running times against MATLAB’s built-in functions is currently not possible, and we only provide comparisons in terms of the number of iterations, number of nonzeros, and the robustness of the new and baseline schemes. The number of iterations and the number of nonzeros provide approximate measures for computational time and memory requirements.
@@ 6 Numerical results
In this section we present the numerical results obtained by our new solver class. Throughout the presented results, we denote by $\star$ if the true residual is greater than the stopping criterion, by $\ddagger$ if the maximum number of iterations is reached, and by §$\S$ if MATLAB runs out of memory. Furthermore, for the baseline (ILUTP) scheme, $\dagger$ indicates a failure in the incomplete factorization, since a zero pivot is encountered. All average values are rounded to one digit after the decimal point.
@@ 6.1 Symmetric case
First, we consider the new scheme without $M$-orthogonalization. For the problems given in Table 1, we present in Table 5 the number of outer GMRES(10) and fGMRES(10) iterations for the baseline and the new schemes, respectively. As expected, the smaller the drop tolerances, the more likely the new method is to converge. The same is true for ILUTP, however, for three cases (reorientation_4, reorientation_8 and tumorAntiAngiogenesis_8) , the final true relative residual norms are below the stopping tolerance even though the preconditioned relative residuals norms indicate that the method has converged. The encounter of a zero pivot is a frequent mode of failure for ILUTP, while the new scheme only failed by reaching the maximum allowed number of iterations without converging to the required stopping tolerance. In Table 6, we report the total number of nonzeros for the new scheme (i.e., $nnz(\widetilde{Z})+nnz(\widetilde{W})$) and ILUTP($\tau$) (i.e., $nnz(L)+nnz(U)$). The best method among those that converge without any failure for each problem is given in bold. The new scheme in all tests achieved fewer nonzeros compared to ILUTP.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
1 17 $\ddagger$ 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$ $\dagger$
2 $\ddagger$ $\ddagger$ 2 $1^{\star}$dagger $\dagger$ $\dagger$ $\dagger$
3 718 $\ddagger$ 2 $1^{\star}$dagger $\dagger$ $\dagger$ $\dagger$
4 14 23 2 2 3 5 $\dagger$ $\dagger$
5 99 $\ddagger$ 4 3 $6^{\star}$ $30^{\star}$dagger $\dagger$
6 $\ddagger$ $\ddagger$ 12 4 7 $\ddagger$ $\dagger$ $\dagger$
7 3 6 2 3 4 6 9 $\dagger$
8 $\ddagger$ $\ddagger$ 1 $3^{\star}$ $3^{\star}$ $7^{\star}$dagger $\dagger$
Table 5: Number of outer fGMRES/GMRES iterations for the symmetric case

### v2

@@ 3.1 The symmetric case
In the symmetric case, we have $B=C$ and $A=A^{T}$, and we compute a sparse approximate nullspace basis, $\widetilde{Z}$, of $B^{T}$. We then assume, see Section 2 for conditions, that ${\mathcal{N}}^{s}=\widetilde{Z}^{T}A\widetilde{Z}$ is positive definite. Since the new scheme involves sparse approximations and is just used as a preconditioner, in the following numerical experiments, we assume that $\widetilde{\mathcal{N}}^{s}$ is positive definite, even if $A$ is indefinite or semidefinite. This allows us to cover a larger number applications than in our initial assumption. The pseudocode of the new scheme for the symmetric case is given in Algorithm 5.
Optionally, we can further $M$-orthogonalize the columns of $\widetilde{Z}$ with respect to $M=A$ using the modified Gram-Schmidt (MGS) procedure given in Algorithm 4. Since it may be too expensive to orthogonalize against all vectors, we also adopt a ’partial’ orthogonalization against a certain number of previous vectors, which we refer to as the window size $w$ (used in line 4) and we employ numerical dropping with a threshold $\tau$ (used in line 7) to ensure sparsity. Although $M$-orthogonalization can be achieved during the oblique conjugation process in Algorithm 2 for the last (nk)(n-k) columns of $V$, we prefer to keep it as a separate stage in our new scheme, since it requires the matrix $M=A$ which may change when solving multiple linear systems with different right-hand sides even though $B$ remains the same.
@@
[…]
@@
[…]
@@
  • Find a particular solution, $\hat{z}_{1}\in\mathbb{R}^{n}$, of $B^{T}z_{1}=t_{2}$ via LSQR
[…]
@@ 3.2 Generalized case
[…]
@@
  • Find a particular solution, $\hat{z}_{1}\in\mathbb{R}^{n}$, of $B^{T}z_{1}=t_{2}$ via LSQR
[…]
@@ 3.3 General case
[…]
@@
5:Solve $\begin{bmatrix}A&B\\ -C^{T}&0\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}=\begin{bmatrix}f\\ g\end{bmatrix}$ via fGMRES, where the systems that involve the preconditioner M $\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}t_{1}\\ t_{2}\end{bmatrix}$ are solved via the nullspace method as follows:
  • Find a particular solution, $\hat{z}_{1}\in\mathbb{R}^{n}$, of $C^{T}z_{1}=t_{2}$ via LSQR
[…]
@@ 4 Test problems
For the generalized case, eight matrices are used (given in Table 2). The first three, namely rajat04, rajat14 and fpga_trans_01, are from the SuiteSparse Matrix Collection. The remaining five test matrices are generated from the IFISS software [11] for the steady-state solution of a linearized 2D Navier-Stokes equations and are obtained from for the lid-driven cavity problem with a spatial discretization based on $Q_{2}/Q_{1}$ elements and a uniform grid, varying the Reynolds number between $100$ and $900$. Since our method is a black-box algebraic approach, we ignore the fact that these problems arise from specific applications and treat them just as given test matrices. The $B$ matrix in this case is numerically rank-deficient. Therefore, in the new scheme, we compute a nullspace basis of size mn+1m-n+1. In Table 2, we provide the smallest and largest eigenvalues of $A+A^{T}$. For three of the test problems in this set, $A+A^{T}$ is positive definite but close to singular, while for the remaining five problems it is robustly positive definite.
For the general case, nine test matrices are used (given in Table 3). Six matrices are obtained from the SuiteSparse Matrix Collection and three are randomly generated. Here, garon1 and garon2 arise in the 2D finite element discretization of the Navier-Stokes equations with inlet and outlet on opposing sides of the geometry. Tolosa matrices (Tols90-4000) arise in the stability analysis of an airplane flight model. The three random matrices are generated as follows: Each block is generated randomly, and to avoid a singular system, a small perturbation of the form $\xi I_{n}$ is added to the $A$ block and $\xi[I_{m}\ 0]^{T}$ to the blocks $C$ and $B$, i.e. for given $m$ and $n$, $A\leftarrow\xi I_{n}+R_{n\times n}$, $B\leftarrow\xi[I_{m}\ 0]^{T}+R_{n\times m}$ and $C^{T}\leftarrow\xi[I_{m}\ 0]+R_{m\times n}$ where $R_{n\times m}$ is a sparse random $n\times m$ matrix with uniformly distributed nonzeros between 0 and $1$, with a density of $1\%$. In our tests, we chose $\xi=10^{-1}$. For all cases, if the right-hand side vector is available, we used it; otherwise, we generated the right-hand side vector from a solution vector of all ones to have a consistent system.
@@
1 reorientation_1 396 281 7,326 $1.7\times 10^{6}$ $1.0\times 10^{9}$
@@
6 goddardRocketProblem_2 434 433 9,058 $1\times 10^{3} 1.1$$1\times 10^{3}$
@@
6 drivencavity_Re500 578 81 10,814 $8.2\times 10^{-4}$ 1.0
7 drivencavity_Re700 578 81 10,814 $2.1\times 10^{-3}$ 1.0
8 drivencavity_Re900 578 81 10,814 $9.1\times 10^{-3}$ 1.0
@@
[…]
@@ 5 Experimental framework
The following numerical results were obtained using MATLAB R2023b on an Apple Macbook Air with an M2 processor and 8 GB of memory. As a baseline of comparison, we use restarted GMRES(m) with $m=10$, preconditioned with MATLAB’s incomplete $LU$ method with pivoting and thresholding, ILUTP($\tau$), where $\tau$ is the drop tolerance. We apply reverse Cuthill-McKee reordering to the systems before computing the incomplete factors by ILUTP($\tau$) to improve the robustness of the baseline method, see [5]. Here, GMRES(m) is preconditioned from the left and right. We note that for the symmetric case, an incomplete $LDL^{T}$ factorization based preconditioner would be more effective and suitable as a baseline, see [16]. Therefore, we use ILUTP($\tau$) for all cases, including the symmetric case. Then in the symmetric case we also use MINRES, preconditioned with the absolute value preconditioner, $L|D|L^{T}$, obtained by modifying the incomplete Bunch-Kaufman-Parlett factorization [16], sym-ildl, where, as in [33], $|D|=V|\Lambda|V^{T}$ if $D$ has a spectral decomposition $V\Lambda V^{T}$.
@@
For our new schemes, we use the restarted and right preconditioned flexible GMRES; fGMRES(m), with $m=10$, as the solver for the outer iteration. We chose fGMRES(m) because the proposed preconditioners (comprised of other iterative solvers) inherently change between iterations. The new schemes have a number of parameters: (a) drop tolerances and thresholds: $\tau_{SAROC},\rho_{SAROC},\tau_{FSAI},\rho_{FSAI}$ and (b) Stopping tolerances for the outer fGMRES, inner LSQR, inner fGMRES and innermost MRS (or inner GC). Furthermore, if $M$-orthogonalization is performed on the nullspace basis, we include the window size ($w_{MGS}$) and drop tolerance ($\tau_{MGS}$) parameters. To illustrate the robustness of the new solver class, we experiment with small, large, and mixed thresholds and tolerances. These are given in Table 4. For ILUTP($\tau$), $\tau$ is set to $10^{-4},10^{-3},10^{-2}$ and $10^{-1}$. For sym-ildl, to compute the incomplete factors, we use the default parameters. That is, the drop tolerance and fill-in are $10^{-3}$ relative to the norm of the column and at most $3$ times the average number of nonzeros per row, respectively. Furthermore, Rook pivoting and AMD reordering are used. We refer the reader to [16] for details of the implementation of these parameters. One of the main difficulties for classical incomplete factorization-based preconditioners is that there is no easy way to predict a good drop tolerance that would work well for a given problem. Sometimes, it is even counterintuitive that decreasing the drop tolerance might result in encountering a zero pivot.
@@
For the new schemes and the baseline methods, we start the iterations with the zero vector and stop when the $L_{2}$ norm of the relative residual drops below $10^{-5}$ or the maximum number (1,000) of iterations is reached. Note that for the case that good starting vectors are available (for example if one has a sequence of slightly changing matrices or right hand sides) then the convergence will be substantially faster.
@@
The residual norms are obtained as a by-product of fGMRES and GMRES. Here, right preconditioned fGMRES produces the true residual norm, and GMRES produces the preconditioned residual norm due to ILUTP-based left and right preconditioning. For these reasons, we also calculate and report the final true relative residual norms. For the new schemes, for the symmetric case, the inner LSQR and CG iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and $\epsilon_{innermost}$ is not used. For the generalized and general case, inner LSQR and fGMRES iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and the innermost MRS iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{innermost}$. Since our new schemes are based on multiple nested loops and external functions, a fair comparison in terms of the running times against MATLAB’s built-in functions is currently not possible, and we only provide comparisons in terms of the number of iterations, number of nonzeros, and the robustness of the new and baseline schemes. The number of iterations and the number of nonzeros provide approximate measures for computational time and memory requirements.
@@ 6 Numerical results
In this section we present the numerical results obtained by our new solver class. Throughout the presented results, we denote by $\star$ if the true residual is greater than the stopping criterion, by $\ddagger$ if the maximum number of iterations is reached, and by $\S$ if MATLAB runs out of memory. Furthermore, for the baseline (ILUTP) scheme, $\dagger$ indicates a failure in the incomplete factorization, since a zero pivot is encountered. All average values are rounded to one digit after the decimal point.
@@ 6.1 Symmetric case
First, we consider the new scheme without $M$-orthogonalization. For the problems given in Table 1, we present in Table 5 the number of outer GMRES(10), MINRES and fGMRES(10) iterations for the baselines and the new scheme, respectively. As expected, the smaller the drop tolerances, the more likely the new method is to converge. The same is true for ILUTP, however, for three cases (reorientation_4, reorientation_8 and tumorAntiAngiogenesis_8), the final true relative residual norms are below the stopping tolerance even though the preconditioned relative residuals norms indicate that the method has converged. The encounter of a zero pivot is a frequent mode of failure for ILUTP, while the new scheme only failed by reaching the maximum allowed number of iterations without converging to the required stopping tolerance. In Table 6, we report the total number of nonzeros for the new scheme (i.e., nnz(Z~)+nnz(W~)nnz($\widetilde${Z})+nnz($\widetilde${W})), ILUTP($\tau$) (i.e., $nnz(L)+nnz(U)$) and sym-ildl (i.e., $nnz(L)+nnz(D)$). The best method among those that converge without any failure for each problem is given in bold. In all tests and in all three cases, the new scheme achieved fewer nonzeros compared to ILUTP and sym-ildl, respectively.
@@
# $\tau_{large}$ $\tau_{mix}$ $\tau_{small}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ sym-ildl
1 17 $\ddagger$ 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$ $\ddagger$
2 $\ddagger$ $\ddagger$ 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 31
3 718 $\ddagger$ 2 $\dagger$ $\dagger$ $\dagger$ $\dagger$ 29
4 14 23 2 3 5 $\dagger$ $\dagger$ 83
5 99 $\ddagger$ 4 $6^{\star}$ $30^{\star}$ $\dagger$ $\dagger$ 117
6 $\ddagger$ $\ddagger$ 12 7 $\ddagger$ $\dagger$ $\dagger$ $\ddagger$
7 3 6 2 4 6 9 $\dagger$ 125
8 $\ddagger$ $\ddagger$ 1 $3^{\star}$ $7^{\star}$ $\dagger$ $\dagger$ 3
Table 5: Number of outer fGMRES/GMRES/MINRES iterations for the symmetric case